points eBook Hint References Check my work Check My Work button is now enabledItem 6 Item 6 0.5 points Exercise 14-11 Bonds; effective interest; adjusting entry [LO14-2] On February 1, 2018, Strauss-Lombardi issued 10% bonds, dated February 1, with a face amount of $930,000. The bonds sold for $855,382 and mature on January 31, 2038 (20 years). The market yield for bonds of similar risk and maturity was 11%. Interest is paid semiannually on July 31 and January 31. Strauss-Lombardi’s fiscal year ends December 31. Required: 1. to 4. Prepare the journal entry to record their issuance by Strauss-Lombardi on February 1, 2018, interest on July 31, 2018 (at the effective rate), adjusting entry to accrue interest on December 31, 2018 and interest on January 31, 2019. (Do not round your intermediate calculations and round your final answers to nearest whole dollar. If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)

  • Record the issuance of the bond on February 1, 2018.

Note: Enter debits before credits.

Date General Journal Debit Credit
February 01, 2018
  • Record the interest on July 31, 2018 (at the effective rate).

Note: Enter debits before credits.

Date General Journal Debit Credit
July 31, 2018
  • Record the accrued interest on December 31, 2018.

Note: Enter debits before credits.

Date General Journal Debit Credit
December 31, 2018
  • Record the interest on January 31, 2019.

Note: Enter debits before credits.

Date General Journal Debit Credit
January 31, 2019
Answer #1
1-Feb Cash 855382
Discount on bonds payable 74618
      Bonds payable 930000
31-Jul Interest expense 47046 =855382*11%/2
      Discount on bonds payable 546
      Cash 46500 =930000*10%/2
31-Dec Interest expense 39230 =(855382+546)*11%/12*5
      Discount on bonds payable 480
       Interest payable 38750 =930000*10%/12*5
31-Jan Interest expense 7846 =(855382+546)*11%/12*1
Interest payable 38750
      Discount on bonds payable 96
      Cash 46500